hhanan:

You invest $500 in an account that has an annual interest rate of 8%, compounded weekly for 12 years. What is the equivalent interest rate and how many times will the money be compounded? How much will you have?

Lulabalula:

grade?

MxxnLight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lulabalula grade? \(\color{#0cbb34}{\text{End of Quote}}\) does it really matter tho- e.e

Lulabalula:

Well i need to know what grade she/he is in so I can see if I can help -_-

XioGonz:

Annual interest rate=8% \[(1+.08/52) ^ 52\] = 1.00153846153846154^52 = 1.0832..... 8.3220......% = interest rate In 12 years it'll will be made up 624 times 12 yr Total = 500 * (1.08322047419671)^12 =1,304.8852611583 = Final Answer 1,304.89

ImBadKidTae:

@14369420